# Quantum computing and the stable set problem

### v1

@@ Abstract
Given an undirected graph, the stable set problem asks to determine the cardinality of the largest subset of pairwise non-adjacent vertices. This value is called the stability number of the graph, and its computation is an NP-hard problem. In this paper, we focus on solving the stable set problem using the D-wave quantum annealer. By formulating the problem as a quadratic unconstrained binary optimization problem with the penalty method, we show that the optimal value of this formulation is equal to the stability number of the graph for certain penalty parameter values. However, D-Wave’s quantum annealer is not an exact solver, so the solutions may be far from the optimum and may not even represent stable sets. To address this, we introduce a post-processing procedure that identifies samples that could lead to improved solutions and extracts stable sets from them. Additionally, we propose a so-called simple CH-partitioning method to handle larger instances that cannot be embedded on D-Wave’s quantum processing unit. Finally, we investigate how different penalty parameter values affect the solutions’ quality. Extensive computational results show that the post-processing procedure significantly improves the solution quality, while the simple CH-partitioning method successfully extends our approach to medium-size instances.
@@ 1.1 Motivation
The stable set problem is a fundamental combinatorial optimization problem. Given an undirected graph, the stable set problem asks to determine the cardinality of the largest subset of pairwise non-adjacent vertices. This value is called the stability number of the graph, and its computation is an NP-complete problem contained in Karp’s list [18] from 1972. Hence, unless P = NP, no polynomial time algorithm can exactly solve the stable set problem in general. However, there are numerous real-life applications of this problem, for instance, in bioinformatics and chemoinformatics [22], coding theory [10], and scheduling [35]. Therefore, efficient algorithms are needed to solve this problem.
@@ 1.2 Quantum annealing
In recent years, a new, although still heuristic, approach for solving combinatorial optimization problems has emerged based on quantum annealing. Quantum annealers leverage quantum mechanics principles to tackle optimization problems. They operate by exploiting one of the quantum phenomena, quantum tunneling, to find the most efficient solution to complex computational challenges. A review of mathematical and theoretical foundations of quantum annealing can be found in [25].
@@
However, challenges persist in achieving fault-tolerant quantum computation. Noise, decoherence, and maintaining quantum coherence remain significant hurdles. D-Wave’s machines are subject to limitations due to factors like thermal fluctuations and imperfections in the qubit layout. As a result, achieving quantum advantage—where quantum computers outperform classical computers for specific tasks—remains an ongoing goal rather than a realized standard in practical applications.
@@ 1.3 Our contribution
    The D-Wave QPU solver could only compute local optimum solutions for instances with less than $150$ vertices. Our results show that very often, these solutions are quite far away from the global solutions in terms of the objective value and may not even be stable sets. To address this, we developed a post-processing procedure that enables us to detect samples that could lead to improved solutions and to extract solutions that are stable sets. Also, we provided theoretical guarantees about the quality of the extracted solutions. The procedure is outlined in Algorithm 1. Table 2 contains detailed numerical results of the solution provided by the QPU solver as part of Algorithm 1.
@@ 1.4 Outline and terminology
The rest of this paper is organized as follows. In Section 2, we give an overview of related work that aims to solve the stable set problem with quantum annealing. A QUBO formulation for the stable set problem is derived in Section 3. In Section 4, we discuss some aspects and shortcomings of the given QUBO formulation and present our post-processing procedure. For larger instances, we propose to use the CH-partitioning method and develop its simplified version in Section 5. In Section 6, we perform extensive computational study and present results for newly proposed approaches. Finally, we conclude with a short discussion of our results in Section 7.
@@
A non-empty graph $G$ is called connected if any of its two vertices are linked by a path in $G$. A maximal connected subgraph of $G$ is called a component of $G$. The complement of a graph $G=(V,E)$ is the graph $\overline{G}=(V,\overline{E})$, where $\overline{E}$ is the set of non-edges in $G$, that is $\overline{E}=\{v_{i}v_{j}\in V\;|\;v_{i}\neq v_{j}\wedge v_{i}v_{j}\notin E\}$. Hence, from the definitions it follows $\alpha(G)=\omega(\overline{G})$ and $\alpha(\overline{G})=\omega(G)$. For a vertex $v_{i}\in V$, we call the set of vertices adjacent to vertex $v_{i}$ the neighborhood set of the vertex $v_{i}$ and denote this set by $N(v_{i})$. For $X\subseteq V$, we denote a subgraph of $G$ induced by $X$ by $G[X]$. The neighborhood of vertices in $X\subseteq V$ is $N(X)=\cup_{v_{i}\in X}N(v_{i})$. The vector of all-ones of length $n$ is denoted by $e_{n}$, and we write $I_{n}$ for the identity matrix of order $n$. If $n$ is clear from the context, we write shortly $e$ and $I$.
@@ 2 Related Work
To solve an optimization problem with D-Wave’s QPU, one should formulate it as an Ising model or a QUBO problem. In [21], respective formulations for many NP-complete and NP-hard problems were introduced. Among others, a QUBO formulation for the stable set was given. The given formulation contains two terms, one for the vertices and one for the edges. While the contribution of a vertex in the solution is rewarded, every edge in the solution is penalized. It is suggested in [21] to set penalization for edge to be strictly greater than a reward for a vertex. Interestingly, in all the related work, authors rewarded a vertex with value one and penalized each edge in the solution with value two.
In this work, we formulate the stable set problem as a QUBO problem, and we demonstrate that an optimal solution can be achieved by assigning a reward of one to a vertex and penalizing an edge with one. Specifically, with an exact QUBO solver, when $\beta\geq\frac{1}{2}$, the optimal value of the QUBO formulation corresponds to the stability number of the graph. However, given the approximate nature of quantum annealing, the solutions obtained may not always be optimal or they may not be stable sets. To address this, we introduce methods for analyzing and post-processing computed solutions to improve them. Our computational analysis suggests that setting the penalization term at $\beta=\frac{1}{2}$ yields better outcomes than the choice of $\beta=1$ used in previous research.
@@ 3.1 Formulation of the stable set problem
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}\forall v_{i}v_{j}\in E,\;x_{i}+x_{j}\leq 1\}.$ (1)
Linear constraints in (1) can be replaced by quadratic ones. The condition that for every edge, at most one end vertex can be in $S$, means that if $v_{i}\in S$, then $x_{j}=0$ for all $v_{i}v_{j}\in E$. Hence, the stability number of a graph $G$ can be computed as the optimal value of the following optimization problem:
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}\forall v_{i}v_{j}\in E,\;x_{i}x_{j}=0\}.$ (2)
The problem (2) can be further reformulated. Let $A$ be the adjacency matrix of a graph $G$. Since $x$ is a binary variable we have that $x_{i}^{2}=x_{i}$, so the edge constraints ${}\forall v_{i}v_{j}\in E,\;x_{i}x_{j}=0$ can be replaced by a single equation $\frac{1}{2}x^{T}Ax=0$. Therefore, the formulation (2) is equivalent to the formulation:
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}\frac{1}{2}x^{T}Ax=0\}.$ (3)
@@
$z^{*}~{}=~{}\min\{f^{T}x+x^{T}Qx\mid x\in\{0,1\}^{n}\}.$
We follow the so-called penalty approach and include the constraint $\frac{1}{2}x^{T}Ax=0$ in the cost function by adding the penalty term $\beta x^{T}Ax$ for a certain value of $\beta$. For the sake of clarity, we define the new cost function as:
@@
Furthermore, for a subset of vertices $X\subset V$, we define:
$\text{cost}(X,\beta)=\text{cost}(x,\beta),$
where $x\in\{0,1\}^{n}$ is the indicator vector for $X$. It is easy to see that we can also write it as:
$\text{cost}(X,\beta)=-|X|+2\beta|E(G[X])|.$
@@
Let $x^{*}\in\{0,1\}^{n}$ be an optimal solution for the problem (3). Since the formulation (3) is equivalent to the formulation (2), we have that $x^{*}_{i}x^{*}_{j}=0$ for all $v_{i}v_{j}\in E$. Hence, the optimal solution $x^{*}$ is a stable set. Furthermore, $x^{*}$ is a feasible solution for the formulation (4), and therefore $\alpha(G)\leq\alpha(G,\beta)$.
Now let $x^{*}\in\{0,1\}^{n}$ be an indicator vector for the optimal subset of vertices $X^{*}$ of (4). If $X^{*}$ is stable set, then $\alpha(G)\geq\alpha(G,\beta)$. Otherwise, there exist $v_{i}v_{j}\in E$, such that $x^{*}_{i}=x^{*}_{j}=1$. If we define new $\tilde{x}$, which differs from $x^{*}$ in the $j$-th component only, which is set to 00, this implies $e^{T}\tilde{x}=e^{T}x^{*}-1$ and $\beta\tilde{x}^{T}A\tilde{x}=\beta(x^{*})^{T}Ax^{*}-2\beta$. Therefore, this change has the following effect on the cost function of (4):
@@
Let $G=(V,E)$ with $V=\{v_{1},v_{2}\}$, $E=\{v_{1}v_{2}\}$ and assume $\beta<\frac{1}{2}$. For $x=(1,1)^{T}$, we have that $\text{cost}(x,\beta)=-e^{T}x+\beta x^{T}Ax=-2+2\beta<-1$. Therefore, the minimum of $\text{cost}(x,\beta)$ over all binary vectors is smaller than $-1$, and thus $\alpha(G,\beta)>1=\alpha(G)$.
Corrolary 3.
Let $G=(V,E)$ be a graph, and let $\tilde{\beta}<\frac{1}{2}$. Then we have
$\alpha(G,\tilde{\beta})\geq\alpha(G),$
and the equality $\alpha(G,\tilde{\beta})=\alpha(G)$ holds if $G$ is a stable set.
@@ 4.1 Analysis of solution
Let $\beta=1$. Then, the penalty component of the QUBO formulation (5) is $x^{T}Ax$. Expanding $x^{T}Ax$, we notice that each quadratic term appears twice, resulting in a penalty of $2$ for each edge. Let us assume there exists a stable set of size $25$ included in a subset of vertices $X\subseteq V$ such that $|X|=30$ and $|E(G[X])|=5$. Then, $\text{cost}(X,1)=-20$, so stable sets of sizes $21,22,23,24$ with respective costs of $-21,-22,-23,-24$ will be more favorable than the set $X$. Therefore, even if only approximate, the solver will likely not return the set $X$ as the optimum. On the other hand, as indicated in Proposition 4, we can easily extract a stable set of size $25$ from the set $X$, meaning that the cost function with $\beta=1$ is not accurately representing the desired solution quality.
From the given example, we note that in the context of the previously given method, choosing the penalization parameter $\beta>\frac{1}{2}$ may discard superior solutions due to the imprecise cost function. In a more general context, for a given set $X\subseteq V$, the value of the cost function is $-|X|+2\beta|E(G[X])|$. However, as established in Proposition 4, we can easily extract a stable set of size $|X|-|E(G[X])|$. Therefore, if we want our cost function to match the size of a stable set that can be extracted from the set $X$, we have to set $12\beta=$ 1. Consequently, $\beta=\frac{1}{2}$ is the most suitable parameter value. If it is any larger, the best candidate set might not be detected, as we have just shown in Example 5, and if it is smaller, Equation (4) does not hold anymore due to Corollary 3.

### v2

@@ Abstract
Given an undirected graph, the stable set problem asks to determine the cardinality of the largest subset of pairwise non-adjacent vertices. This value is called the stability number of the graph, and its computation is an NP-hard problem. In this paper, we solve the stable set problem using the D-wave quantum annealer. By formulating the problem as a quadratic unconstrained binary optimization problem with the penalty method, we show its optimal value equals the graph’s stability number for specific penalty values. However, D-Wave’s quantum annealer is a heuristic, so the solutions may be far from the optimum and may not represent stable sets. To address these, we introduce a post-processing procedure that identifies samples that could lead to improved solutions. Additionally, we propose a partitioning method to handle larger instances that cannot be embedded on D-Wave’s quantum processing unit. Finally, we investigate how different penalty parameter values affect the solutions’ quality. Extensive computational results show that the post-processing procedure significantly improves the solution quality, while the partitioning method successfully extends our approach to medium-size instances.
@@ 1.1 Motivation
The stable set problem is a fundamental combinatorial optimization problem. Given an undirected simple graph $G$, the stable set problem asks for the largest subset of vertices that are pairwise non-adjacent, known as the stability number of the graph. It is well-known that stable sets in $G$ correspond to cliques in the complement graph $\overline{G}$, making the stable set problem in $G$ equivalent to the maximum clique problem in $\overline{G}$. The maximum clique problem, and thus the stable set problem, is an NP-hard problem included in Karp’s 1972 list of NP-complete problems [18]. Hence, unless P = NP, no polynomial time algorithm can exactly solve the stable set problem in general. However, there are numerous real-life applications of this problem, for instance, in bioinformatics and chemoinformatics [22], coding theory [10], and scheduling [35]. Therefore, efficient algorithms are needed to solve this problem.
@@ 1.2 Quantum annealing
In recent years, a new heuristic approach for solving combinatorial optimization problems has emerged based on quantum annealing. Quantum annealers leverage quantum mechanics principles to tackle optimization problems. They operate by exploiting one of the quantum phenomena, quantum tunneling, to find the most efficient solution to complex computational challenges. A review of mathematical and theoretical foundations of quantum annealing can be found in [25].
@@
However, challenges persist in achieving fault-tolerant quantum computation. Noise and maintaining quantum coherence remain significant hurdles. D-Wave’s machines are subject to limitations due to factors like thermal fluctuations and imperfections in the qubit layout. As a result, achieving quantum advantage—where quantum computers outperform classical computers for specific tasks—remains an ongoing goal rather than a realized standard in practical applications.
@@ 1.3 Our contribution
    The D-Wave QPU solver could only compute local optimum solutions for instances with less than $150$ vertices. Our results show that very often, these solutions are quite far away from the global solutions in terms of the objective value and may not even be stable sets. To address these, we developed a post-processing procedure that enables us to detect samples that could lead to improved solutions and to extract solutions that are stable sets. Also, we provide theoretical guarantees about the quality of the extracted solutions. The procedure is outlined in Algorithm 1. Table 2 contains detailed numerical results of the solution provided by the QPU solver as part of Algorithm 1.
@@ 1.4 Outline and terminology
The rest of this paper is organized as follows. In Section 2, we give an overview of related work that aims to solve the stable set problem with quantum annealing. A QUBO formulation for the stable set problem is derived in Section 3. In Section 4, we discuss some aspects and shortcomings of the given QUBO formulation and present our post-processing procedure. For larger instances, we propose to use the CH-partitioning method and develop its modified version in Section 5. In Section 6, we perform extensive computational study and present results for newly proposed approaches. Finally, we conclude with a short discussion of our results in Section 7.
@@
A non-empty graph $G$ is called connected if any of its two vertices are linked by a path in $G$. A maximal connected subgraph of $G$ is called a component of $G$. The complement of a graph $G=(V,E)$ is the graph $\overline{G}=(V,\overline{E})$, where $\overline{E}$ is the set of non-edges in $G$, that is $\overline{E}=\{v_{i},v_{j}\in V\;|\;v_{i}\neq v_{j}\wedge\{v_{i},v_{j}\}\notin E\}$. Hence, from the definitions it follows $\alpha(G)=\omega(\overline{G})$ and $\alpha(\overline{G})=\omega(G)$. The relationship between the stable set and the maximum clique problem is visualized in Figure 1.
$v_{5}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$
$v_{5}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$
Figure 1: The left figure shows a graph $G$ on five vertices. In this graph, the following subsets of vertices are stable sets: $\{v_{1},v_{2}\}$, $\{v_{1},v_{3}\}$, $\{v_{1},v_{5}\}$, $\{v_{1},v_{3},v_{5}\}$, $\{v_{3},v_{5}\}$, $\{v_{4},v_{5}\}$. The right figure shows the complement graph $\overline{G}$, where these same subsets form cliques. In particular, the maximum stable set $\{v_{1},v_{3},v_{5}\}$ in $G$ is equivalent to the maximum clique in $\overline{G}$, and therefore, $\alpha(G)=\omega(\overline{G})=3$.
For a vertex $v_{i}\in V$, the set of vertices adjacent to $v_{i}$ is called its neighborhood. The closed neighborhood of vertex $v_{i}$ includes the vertex $v_{i}$ and its neighborhood. We denote this closed neighborhood set by $N(v_{i})$. For $X\subseteq V$, we denote a subgraph of $G$ induced by $X$ by $G[X]$. The closed neighborhood of vertices in $X\subseteq V$ is $N(X)=\cup_{v_{i}\in X}N(v_{i})$. The vector of all-ones of length $n$ is denoted by $e_{n}$, and we write $I_{n}$ for the identity matrix of order $n$. If $n$ is clear from the context, we write shortly $e$ and $I$.
@@ 2 Related Work
To solve an optimization problem with D-Wave’s QPU, one should formulate it as an Ising model or a QUBO problem. In [21], respective formulations for many NP-complete and NP-hard problems were introduced. Among others, a QUBO formulation for the stable set was given. The given formulation contains two terms, one for the vertices and one for the edges. While the contribution of a vertex in the solution is rewarded, every edge in the solution is penalized. It is suggested in [21] to set penalization for edge to be strictly greater than a reward for a vertex.
In the related work discussed in this section, the authors focused on penalizing edges without rewarding vertices. Specifically, when adapting their formulations to the QUBO formulation given in this work—where vertices are not rewarded, but edges are penalized using a penalty parameter $\beta$—we observe that a value of $\beta=1$ was used in all these studies.
In contrast, we demonstrate that, with an exact QUBO solver, the optimal value of the given QUBO formulation corresponds to the stability number of the graph when the penalty term $\beta$ is set to at least $\frac{1}{2}$. However, given the approximate nature of quantum annealing, the solutions obtained may not always be optimal or they may not be stable sets. To address these issues, we introduce methods for analyzing and post-processing computed solutions to improve them. Our computational analysis suggests that setting the penalization term at $\beta=\frac{1}{2}$ yields better outcomes than the choice of $\beta=1$ used in previous research.
@@ 3.1 Formulation of the stable set problem
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}\forall\{v_{i},v_{j}\}\in E,\;x_{i}+x_{j}\leq 1\}.$ (1)
Linear constraints in (1) can be replaced by quadratic ones. The condition that for every edge, at most one end vertex can be in $S$, means that if $v_{i}\in S$, then $x_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the stability number of a graph $G$ can be computed as the optimal value of the following optimization problem:
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}\forall\{v_{i},v_{j}\}\in E,\;x_{i}x_{j}=0\}.$ (2)
The problem (2) can be further reformulated. Let $A$ be the adjacency matrix of a graph $G$. Since $x$ is a binary variable we have that $x_{i}^{2}=x_{i}$, so the edge constraints {${}\forall\{v_{i},v_{j}\}\in E,\;x_{i}x_{j}=0$ can be replaced by a single equation $\frac{1}{2}x^{T}Ax=0$. Therefore, the formulation (2) is equivalent to the formulation:
$\alpha(G)~{}=~{}\max\{e^{T}x\mid~{}x\in\{0,1\}^{n}\wedge~{}\frac{1}{2}x^{T}Ax=0\}.$
@@
$z^{*}~{}=~{}\min\{f^{T}x+x^{T}Qx\mid x\in\{0,1\}^{n}\}.$ (3)
We follow the so-called penalty approach and penalize edge constraints. Specifically, since $\frac{1}{2}x^{T}Ax=0~{}\Leftrightarrow~{}x^{T}Ax=0$, we penalize the term $x^{T}Ax$ in the cost function by adding the penalty term $\beta x^{T}Ax$ for a certain value of $\beta\geq 0$. For the sake of clarity, we define the new cost function as:
@@
Furthermore, for a subset of vertices $X\subseteq V$, we define:
$\text{cost}(X,\beta)=\text{cost}(x,\beta),$ (4)
where $x\in\{0,1\}^{n}$ is the indicator vector for $X$. The role of the penalty parameter $\beta$ can be understood as follows.
Observation 1.
In the given cost function (4), if there exists an edge between vertices $v_{i}$ and $v_{j}$, and if $x_{i}=x_{j}=1$, the term $x^{T}Ax$ counts this edge twice. Consequently, for $\beta\geq 0$, the penalty term $\beta x^{T}Ax$ penalizes each edge in $G[X]$ with the value of $2\beta$.
Based on Observation 1, we can rewrite the cost function (4) as:
$\text{cost}(X,\beta)=-|X|+2\beta|E(G[X])|.$ (5)
@@
Let $x^{*}\in\{0,1\}^{n}$ be an optimal solution for the problem (3). Since the formulation (3) is equivalent to the formulation (2), we have that $x^{*}_{i}x^{*}_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the optimal solution $x^{*}$ is a stable set. Furthermore, $x^{*}$ is a feasible solution for the formulation (6), and therefore $\alpha(G)\leq\alpha(G,\beta)$.
Now let $x^{*}\in\{0,1\}^{n}$ be an indicator vector for an optimal subset of vertices $X^{*}$ of (6). If $X^{*}$ is stable set, then $\alpha(G)\geq\alpha(G,\beta)$. Otherwise, there exist $\{v_{i},v_{j}\}\in E$, such that $x^{*}_{i}=x^{*}_{j}=1$. If we define new $\tilde{x}$, which differs from $x^{*}$ in the $j$-th component only, which is set to 00, this implies $e^{T}\tilde{x}=e^{T}x^{*}-1$ and $\beta\tilde{x}^{T}A\tilde{x}=\beta(x^{*})^{T}Ax^{*}-2\beta$. Therefore, this change has the following effect on the cost function of (6):
@@
Let $G=(V,E)$ with $V=\{v_{1},v_{2}\}$, $E=\{\{v_{1},v_{2}\}\}$ and assume $\beta<\frac{1}{2}$. For $x=(1,1)^{T}$, we have that $\text{cost}(x,\beta)=-e^{T}x+\beta x^{T}Ax=-2+2\beta<-1$. Therefore, the minimum of $\text{cost}(x,\beta)$ over all binary vectors is smaller than $-1$, and thus $\alpha(G,\beta)>1=\alpha(G)$.
Corollary 4.
Let $G=(V,E)$ be a graph, and let $\beta<\frac{1}{2}$. Then we have
$\alpha(G,\beta)\geq\alpha(G),$
and the equality $\alpha(G,\beta)=\alpha(G)$ holds if $G$ is a stable set.
Proof.
The inequality $\alpha(G,\beta)\geq\alpha(G)$ follows from the proof of Lemma 2. We now show the equality. Assume $G$ is a stable set. Then $\alpha(G)=|V|$, the total number of vertices. Furthermore, for a stable set $G$, the edge set $E(G[X])$ is empty for any subset $X\subseteq V$, so $\text{cost}(X,\beta)=-|X|$. Thus, the minimum cost function over all subsets $X$ is achieved when $X=V$, yielding
$\alpha(G,\beta)=-\min\{-|X|\mid X\subseteq V\}=|V|.$
Therefore, $\alpha(G,\beta)=|V|=\alpha(G)$. ∎
@@ 4.1 Analysis of solution
Let $\beta=1$. Then, the penalty component of the QUBO formulation (7) is $x^{T}Ax$. Expanding $x^{T}Ax$, we notice that each quadratic term appears twice according to Observation 1, resulting in a penalty of $2$ for each edge. Let us assume there exists a stable set of size at least of $25$ included in a subset of vertices $X\subseteq V$ such that $|X|=30$ and $|E(G[X])|=5$. Then, $\text{cost}(X,1)=-20$, so stable sets of sizes $21,22,23,24$ with respective costs of $-21,-22,-23,-24$ will be more favorable than the set $X$. Therefore, even if only approximate, the solver will likely not return the set $X$ as the optimum. On the other hand, as indicated in Proposition 5, we can easily extract a stable set of size $25$ from the set $X$, meaning that the cost function with $\beta=1$ is not accurately representing the desired solution quality.
From the given example, we note that in the context of the previously given method, choosing the penalization parameter $\beta>\frac{1}{2}$ may discard superior solutions due to the imprecise cost function. To explain this in a more general context, we first need to prove the following statement.
Lemma 7.
For every $\epsilon>0$ there exist a graph $G$ and a subset $X\subseteq V$ such that
$1\displaystyle 2\epsilon|E(G[X])|>$ 1 . (8)
Proof.
We will prove the lemma by providing an example of such graph. Let $\varepsilon>0$ and set $m=\lceil\frac{1}{2\epsilon}+1\rceil$. The graph which will serve as an example is a union of $m$ edges, i.e, we have $|E|=m$, $|V|=2m$ and $\deg(v)=1$ for every $v\in V$. For $X=V$ it follows that:
$|E(G[X])|=|E(G[V])|=|E|=m>\frac{1}{2\epsilon},$
so the inequality (8) holds for this graph. ∎
Now, suppose we have a graph $G=(V,E)$, a set $X\subseteq V$ and $\beta=\frac{1}{2}+\epsilon$ where $\epsilon>0$. The value of the cost function (5) is:
$-|X|+2\beta|E(G[X])|=-(|X|-(1+2\epsilon)|E(G[X])|).$
To improve clarity and simplify the explanation, we will flip the sign of the cost function and approach the problem as a maximization problem. This adjustment transforms the cost function into:
$cost=|X|-(1+2\epsilon)|E(G[X])|.$
As established in Proposition 5, we can easily extract a stable set of size:
$solution=|X|-|E(G[X])|$
out of the $X$. Since we are dealing with an arbitrary graph, we can select $G$ and the set $X\subseteq V$ such that:
$solution-cost=2\epsilon|E(G[X])|>1$
holds, which implies that $solution-1>cost$. Existence of such graph is ensured by Lemma 7. Since the cost function value of a stable set is equal to the cardinality of the stable set itself, this means that some stable set $X^{\prime}$ of cardinality $|X^{\prime}|=solution-1$ has better cost function value than our set $X$, even though we can easily extract a stable set of size $solution$ out of the set $X$. Therefore, the cost function imprecisely evaluates our solution and can potentially favor some inferior solutions. To avoid this issue, $\epsilon$ should be set to 00, ensuring that $solution-cost$ cannot exceed $1$. Consequently, $\beta=\frac{1}{2}$ is the most suitable parameter value. If it is any larger, the best candidate set might not be detected as we have just shown, and if it is smaller, Equation (6) does not hold anymore due to Corollary 4.